# Extended Landauer-Büttiker Formula for Current through Open Quantum Systems with Gain or Loss

### v3

@@ Abstract
Introduction.— Studying quantum transport properties is essential for understanding condensed-matter systems and has broad applications, such as in various electronic devices, atomic and photonic technologies, nanotechnology, and quantum computing review0 ; review1 ; review2 ; review3 ; review4 ; review5 ; review6 . One of the most influential frameworks for understanding transport is the Landauer-Büttiker formula LB1 ; LB2 , which expresses current in terms of the system’s transmission coefficient and the electron distribution within connected leads. Beyond electrons, this formalism applies universally to any quantum system–including atomic BLtransport1 ; BLtransport2 ; BLtransport3 ; BLtransport4 ; BLtransport5 , photonic BLtransport6 ; BLtransport7 , and phononic BLtransport8 systems–where transport reduces to a scattering problem with well-defined transmission channels BLnote . The system described by this formula, aside from being coupled to two leads, is not influenced by any form of dissipation. However, dissipation is widespread, and understanding its effects on quantum transport is a fundamentally important issue.
@@
Figure 2: (a) $J^{0}$ as a function of $\gamma_{g}/\gamma_{l}$ and $\mu/k_{B}T$, in units of $e^{2}/h$, with $k_{B}T=20$, $\gamma_{l}=0.2$, and the system size $N=40$. (b1) Scheme of a two-site system with local monitoring. (b2) $J^{0}$ as a function of monitoring strength $\gamma_{0}$. The blue and red curves represent monitoring at the first and second sites, respectively, while the green curve represents monitoring at both sites. (c) $J^{0}$ as a function of disorder strength $V$ with 200200 samples and $N=100$. The black dashed line represents the envelope curve of the current. (d) The sample average of $(J^{0})^{2}$. The red dashed line is the fitting function $\langle(J^{0})^{2}\rangle\sim V^{1.91}$. Other parameters are (b-d) $\mu=0.1$, $k_{B}T=10^{-4}$, $\gamma_{l}=0.1$, and (a-d) $t_{S}=1$, $(\bm{\Gamma}_{L})_{11}=(\bm{\Gamma}_{R})_{NN}=1.1$.
@@ Acknowledgements.
IV I. Comparative analysis: Markovian reservoirs and conventional leads
@@
V II. Derivation of the Lindblad master equation
@@
where $\lambda_{jk}^{(m)}$ denotes the coupling strength. It is assumed that the reservoirs satisfy the Markovian condition, meaning the relaxation time $\tau_{B}$ of reservoirs is significantly shorter than the dynamical evolution timescale $\Delta t$ of the central systems Citro ; Brasil .
@@
Over a time interval $\Delta t$, the change in the density matrix $\Delta\bar{\rho}(t)=\bar{\rho}(t+\Delta t)-\bar{\rho}(t)$ can be calculated by
$\Delta\bar{\rho}(t)=-i\int_{t}^{t+\Delta t}dt^{{}^{\prime}}[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),\bar{\rho}(t^{{}^{\prime}})]-\int_{t}^{t+\Delta t}dt^{{}^{\prime}}\int_{t}^{t^{{}^{\prime}}}dt^{{}^{\prime\prime}}[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),[\sum_{n}\bar{H}_{SB}^{(n)}(t^{{}^{\prime\prime}}),\bar{\rho}(t^{{}^{\prime\prime}})]].$ (S6)
Assuming the central system and the reservoirs are weak correlated and the reservoirs are in equilibrium, such that then density matrix can be decomposed by $\bar{\rho}(t)=\bar{\rho}_{S}(t)\otimes\prod_{m}\rho_{B}^{(m)}$. Substituting it into Eq. (S6) and tracing over the reservoirs gives
$\Delta\bar{\rho}_{S}(t)=-\int_{t}^{t+\Delta t}dt^{{}^{\prime}}\int_{t}^{t^{{}^{\prime}}}dt^{{}^{\prime\prime}}Tr_{B}\bigg[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),[\sum_{n}\bar{H}_{SB}^{(n)}(t^{{}^{\prime\prime}}),\bar{\rho}_{S}(t)\otimes\rho_{B}]\bigg].$ (S7)
Due to the coupling between the system and the reservoir is linear, the term $\bar{H}_{SB}^{(m)}(t)$ contains odd number of operators $a^{(m)}$ and $a^{(m)\dagger}$, hence $Tr_{B}[\bar{H}_{SB}^{(m)}(t),\bar{\rho}(t)]=0$. Furthermore, the coarse-grained approximation $\bar{\rho}(t^{{}^{\prime\prime}})\approx\bar{\rho}(t)$ is applied, as $\bar{\rho}_{S}$ evolves much slower than the reservoirs. Substituting Eq. (S4) into Eq. (S7) we derive
@@
V.1 III. Definition of particle and energy currents
@@
V.2 IV. Keldysh theory of open fermionic systems
@@
V.3 V. Derivation details of extended Landauer-Büttiker formula
@@
VI VI. Steady-state currents in bosonic systems with gain and loss
@@
VII VII. Effect of inversion symmetry in gain or loss terms on the current: a two-site example
@@
VIII VIII. Disorder-induced current generation
@@
IX IX. Response current and the Wiedemann-Franz law
@@
X X. $J_{0}^{0}$ variation with size in the presence of skin effect

### v4

@@ Abstract
Introduction.— Studying quantum transport properties is essential for understanding condensed-matter systems and has broad applications, such as in various electronic devices, atomic and photonic technologies, nanotechnology, and quantum computing review0 ; review1 ; review2 ; review3 ; review4 ; review5 ; review6 . One of the most influential frameworks for understanding transport is the Landauer-Büttiker formula LB1 ; LB2 , which expresses current in terms of the system’s transmission coefficient and the electron distribution within connected leads. Beyond electrons, this formalism applies universally to any quantum system–including atomic BLtransport1 ; BLtransport2 ; BLtransport3 ; BLtransport4 ; BLtransport5 , photonic BLtransport6 ; BLtransport7 , and phononic BLtransport8 systems–where transport reduces to a scattering problem with well-defined transmission channels BLnote . The system described by this formula, in its original coherent formulation, is not influenced by any form of dissipation other than the coupling to the two leads explainOLB ; ButtikerVP . However, dissipation is widespread, and understanding its effects on quantum transport is a fundamentally important issue.
@@
Figure 2: (a) $J^{0}$ as a function of $\gamma_{g}/\gamma_{l}$ and $\mu/k_{B}T$, in units of $1/h$, with $k_{B}T=20$, $\gamma_{l}=0.2$, and the system size $N=40$. (b1) Scheme of a two-site system with local monitoring. (b2) $J^{0}$ as a function of monitoring strength $\gamma_{0}$. The blue and red curves represent monitoring at the first and second sites, respectively, while the green curve represents monitoring at both sites. (c) $J^{0}$ as a function of disorder strength $V$ with 200200 samples and $N=100$. The black dashed line represents the envelope curve of the current. (d) The sample average of $(J^{0})^{2}$. The red dashed line is the fitting function $\langle(J^{0})^{2}\rangle\sim V^{1.91}$. Other parameters are (b-d) $\mu=0.1$, $k_{B}T=10^{-4}$, $\gamma_{l}=0.1$, and (a-d) $t_{S}=1$, $(\bm{\Gamma}_{L})_{11}=(\bm{\Gamma}_{R})_{NN}=1.1$.
@@ Acknowledgements.
II I. Comparative analysis: Markovian reservoirs and conventional leads
@@
III II. Derivation of the Lindblad master equation
@@
where $\lambda_{jk}^{(m)}$ denotes the coupling strength.
@@
Over a time interval $\Delta t$, the change in the density matrix is given by
$\Delta\bar{\rho}(t)\equiv\bar{\rho}(t+\Delta t)-\bar{\rho}(t)=-i\int_{t}^{t+\Delta t}dt^{{}^{\prime}}[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),\bar{\rho}(t^{{}^{\prime}})].$ (S6)
Eq. (S6) can be iterated by inserting for $\bar{\rho}(t^{{}^{\prime}})$ again, which leads to
$\Delta\bar{\rho}(t)=-i\int_{t}^{t+\Delta t}dt^{{}^{\prime}}[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),\bar{\rho}(t)]-\int_{t}^{t+\Delta t}dt^{{}^{\prime}}\int_{t}^{t^{{}^{\prime}}}dt^{{}^{\prime\prime}}[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),[\sum_{n}\bar{H}_{SB}^{(n)}(t^{{}^{\prime\prime}}),\bar{\rho}(t^{{}^{\prime\prime}})]].$ (S7)
Assuming that the reservoirs satisfy the Markovian condition, meaning the relaxation time $\tau_{B}$ of reservoirs is signifcantly shorter than the dynamical evolution timescale $\Delta t$ of the central systems, one can prove that the correlations between system and environment can be neglected Citro ; Brasil . Then the total density operator can be factorized as (Born approximation)
$\bar{\rho}(t)\approx\bar{\rho}_{S}(t)\otimes\rho_{B},\quad\rho_{B}=\prod_{m}\rho_{B}^{(m)}.$ (S8)
Under the Markov coarse-graining condition, the reduced density matrix varies slowly within the integration window, allowing the replacement $\bar{\rho}(t^{{}^{\prime}})\approx\bar{\rho}(t^{{}^{\prime\prime}})\approx\bar{\rho}(t)$. Substituting this factorization Eq. (S8) into Eq. (S7) and tracing over the reservoir degrees of freedom yields
$\Delta\bar{\rho}_{S}(t)\equiv\bar{\rho}_{S}(t+\Delta t)-\bar{\rho}_{S}(t)=\mathrm{Tr}_{B}\left[\Delta\bar{\rho}(t)\right]=-\int_{t}^{t+\Delta t}dt^{{}^{\prime}}\int_{t}^{t^{{}^{\prime}}}dt^{{}^{\prime\prime}}Tr_{B}\bigg[\sum_{m}\bar{H}_{SB}^{(m)}(t^{{}^{\prime}}),[\sum_{n}\bar{H}_{SB}^{(n)}(t^{{}^{\prime\prime}}),\bar{\rho}_{S}(t)\otimes\rho_{B}]\bigg].$ (S9)
Because the system-reservoir coupling is linear, the term $\bar{H}_{SB}^{(m)}(t)$ contains a single (odd number of) bath creation/annihilation operator(s). For an equilibrium reservoir $Tr_{B}[a^{(m)}\rho_{B}]=Tr_{B}[a^{(m)\dagger}\rho_{B}]=0$, so the first-order term vanishes and $Tr_{B}[\bar{H}_{SB}^{(m)}(t),\bar{\rho}(t)]=0$. Substituting Eq. (S4) into Eq. (S9) we derive
@@
III.1 III. Definition of particle and energy currents
@@
III.2 IV. Keldysh theory of open fermionic systems
@@
III.3 V. Derivation details of extended Landauer-Büttiker formula
@@
IV VI. Steady-state currents in bosonic systems with gain and loss
@@
V VII. Effect of inversion symmetry in gain or loss terms on the current: a two-site example
@@
VI VIII. Disorder-induced current generation
@@
VII IX. Response current and the Wiedemann-Franz law
@@
VIII X. $J_{0}^{0}$ variation with size in the presence of skin effect